OpenStudy (anonymous):

two number cubes are rolled. what is the probability that the sum will be even ?

OpenStudy (sandra):

ok so the probability of something occurring is always (#cases where event occurs)/(#total number of possible cases)

OpenStudy (sandra):

so in the case of dice, there are 36 possible roles of the dice (6 possiblities for the first , and 6 possibilities for the second - so 6 x 6)

OpenStudy (anonymous):

alright. which is 36?

OpenStudy (sandra):

so by brute force, you can actually list all of the combinations, and then count the number that are even: Dice 1, Dice 2, Sum, Odd/even 1. 1, 1, 1+1=2, even 2. 1, 2, 1+2=3, odd 3. 1, 3, 1+3=4, even and so on... until you've listed all 36

OpenStudy (anonymous):

wait (6) (6)^2?

OpenStudy (sandra):

and then the answer is the number of even possibilities/36

OpenStudy (sandra):

you were correct - there are 36 possible combinations when you roll a dice

OpenStudy (anonymous):

so 36 possible even number? 6/36?

OpenStudy (sandra):

1 and 1, 1 and 2, 1 and 3, and so on. But since you know that the number of possiblities for each individual dice is 6, the total number of possibilities when you roll two of them is 6 x 6 = 36

OpenStudy (sandra):

now you want to know, out of those 36 possible combinations, how many add up to be an even number?

OpenStudy (sandra):

you can brute force the solution by listing them all, and their sums, and know that the probability is #evensums/36

OpenStudy (anonymous):

yea. probability some will be even

OpenStudy (anonymous):

(6/36) (6/36)

OpenStudy (sandra):

hmm, the question said "sum", not "some"

OpenStudy (anonymous):

sum****** my bad lol

OpenStudy (sandra):

so a sum is even when you add an even and an even, OR an odd and an odd

OpenStudy (anonymous):

when you add two even numbers together you get an even sum and when you add two odd nubers together you gat an even sum otherwise it is odd.

OpenStudy (sandra):

great minds think alike =)

OpenStudy (sandra):

so what's the probability that both are even numbers? Well, since there's 3/6 chance for the first dice being even, and 3/6 chance the second dice is even, there's a 3/6 * 3/6 = 1/4 chance of both being even

OpenStudy (sandra):

what's the probability of them both being odd?